OpenStudy (animalover02):

I WILL METAL!!!!!!!!!!!!! Given 4x − 8y = 8, transform the equation into slope-intercept form.

OpenStudy (animalover02):

@563blackghost @Will.H

OpenStudy (caleb4324):

4x - 8y = 8 Subtract 4x from each side. -8y = 8 - 4x Divide -8 to each. y = -1 + 1/2x So, the answer is y = 1/2x - 1

OpenStudy (animalover02):

@fixx

OpenStudy (fixxer):

do you even try? or do you copy question without even reading it?

OpenStudy (animalover02):

i have a bunch but i have little time so im needing help:)

OpenStudy (fixxer):

I see, but on the exam you cant get help. Getting help is ok, but when you got answeres to so many questions just like this it dosent seems like you even glance at the answere you get